Cell Phone Use and Brain Cancer

The association between cellular telephone use and the risk of brain cancer was investigated in a case-control study. The study included 475 cases and 400 controls and the following results were seen:

Cases   Controls

Cellular   Yes       270   200

Phone

User       No       205   200

       Total       475   400  

Calculate the odds ratio based on these data. 7 points.

Calculate the 95% confidence interval for the OR you just calculated. 7 points.

Interpret the 95% confidence interval. 7 points.
Answer #1

Answer:

Cell Phone Use and Brain Cancer

The association between cellular telephone use and the risk of brain cancer was investigated in a case-control study. The study included 475 cases and 400 controls and the following results were seen:

Cases   Controls

Cellular   Yes       270   200

Phone

User       No       205   200

       Total       475   400  

Calculate the odds ratio based on these data. 7 points.

Odds ratio =ad/bc

= (270*200)/(205*200)

=1.3171

Calculate the 95% confidence interval for the OR you just calculated. 7 points.

se( log(OR)) = sqrt( 1/270+1/200+1/205+1/200) =0.1363

95% CI for log(OR) = (log(OR)-1.96* se( log(OR)), log(OR)+1.96* se( log(OR)))

= (0.2754-1.96*0.1363, 0.2754-1.96*0.1363)

=(0.0083, 0.5425)

95% CI for OR = ( exp(0.0083), exp(0.5425) )

= (1.0083, 1.7203)

Interpret the 95% confidence interval. 7 points.

We are 95% confident that true OR falls in the interval (1.0083, 1.7203).